OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4).

mathslover (mathslover):

Just put the value of x = 4 , in g(x) \[\huge{g(4)=4^2+3}\] can u solve this ?

OpenStudy (anonymous):

replace x in g(x) by 4

OpenStudy (anonymous):

substitute x = 4 (4x2) + 3 = 11